OpenStudy (anonymous):

A pizza place offers 10 topping choices for their pizzas. How many different two-topping pizzas can you order? A.2 B.45 C.90 D.100

OpenStudy (anonymous):

\(\frac{10\times 9}{2}\)

OpenStudy (anonymous):

ten for the first topping, nine for the second, and since order does not count you divide by 2

OpenStudy (anonymous):

i should say "ten choices for the first topping, then nine for the second, so by the counting principle it is \(10\times 9\) and since order does not matter, i.e. choosing anchovies and sausage is the same as choosing sausage and anchovies, you divide by 2